OpenStudy (wwe123):

quadratic equation..... if a≠b and a²=5a -3, b²=5b-3,then that equation whose roots are...

OpenStudy (wwe123):

\[\frac{ a }{ b } and \frac{ b }{ a }\]

OpenStudy (wwe123):

hint. a,b are the roots of the equation x²=5x-3

OpenStudy (anonymous):

yes, that is for sure so you know what they are

OpenStudy (anonymous):

on is \(\frac{5+\sqrt{3}}{2}\) and the other is its conjugate

OpenStudy (anonymous):

therefore \[\frac{a}{b}=\frac{5+\sqrt{3}}{5-\sqrt{3}}\]

OpenStudy (anonymous):

or its reciprocal, either way so you can find the equation by knowing the roots, although somehow i get the feeling there is a slicker way to do this

OpenStudy (wwe123):

how u get \[\frac{ 5+\sqrt{3} }{2 }\] ??

OpenStudy (wwe123):

i got u & also done it

OpenStudy (wwe123):

thanks @satellite73